# Optimal Error Analysis of Channel Estimation for IRS-assisted MIMO Systems

### v1

@@ Discussion
IV Nonconvex Optimization Algorithms
In this section, our focus will be on designing algorithms for estimating the channel matrices within the framework of the TT-based ToT regression model. Initially, we will concentrate on the single-hop IRS-assisted MIMO system and subsequently extend our approach to the multi-hop IRS-assisted MIMO system. Note that, given that the following algorithms can be readily extended to handle the unknown $\mathcal{S}$ scenario, our attention is directed towards the known $\mathcal{S}$ case at the receiver.
Single-hop IRS-assisted MIMO system
$(i)$ One common assumption is that the pilot signal matrix $\bm{X}$ is a truncated discrete Fourier transform (DFT) matrix or a semi-unitary matrix, as discussed in previous works [13, 31]. However, we can extend this assumption by considering the existence of the pseudoinverse for $\bm{X}$. Expanding upon the equivalent form of the TT format $\mathcal{B}=[\bm{H},\mathcal{S},\bm{G}]$ [76], denoted as $\mathcal{B}^{\langle 1\rangle}=\bm{H}R(\mathcal{S})(\bm{G}\otimes{\bf I}_{K})$ and $\mathcal{B}^{\langle 2\rangle}=({\bf I}_{K}\otimes\bm{H})L(\mathcal{S})\bm{G}$ (The expressions for $L(\mathcal{S})$ and $R(\mathcal{S})$ are provided in (31) and (32) of Appendix A.)  [^6] Let $\mathcal{X}^{\langle i\rangle}\in\mathbb{C}^{(d_{1}\cdots d_{i})\times(d_{i+1}\cdots d_{N})}$ denote the $i$-th unfolding matrix of the tensor $\mathcal{X}\in\mathbb{C}^{d_{1}\times\dots\times d_{N}}$. The $(s_{1}\cdots s_{i},s_{i+1}\cdots s_{N})$-th element of $\mathcal{X}^{\langle i\rangle}$ is given by $\mathcal{X}^{\langle i\rangle}(s_{1}\cdots s_{i},s_{i+1}\cdots s_{N})=\mathcal{X}(s_{1},\dots,s_{N})$. Here, $s_{1}\cdots s_{i}$ and $s_{i+1}\cdots s_{N}$ represent the row and column indices, respectively, where the row index is calculated as $s_{1}+d_{1}(s_{2}-1)+\cdots+d_{1}\cdots d_{i-1}(s_{i}-1)$ and the column index as $s_{i+1}+d_{i+1}(s_{i+2}-1)+\cdots+d_{i+1}\cdots d_{N-1}(s_{N}-1)$., we can formulate the following loss function:
$\min_{\widehat{\bm{H}}\in\mathbb{C}^{LP\times N}}\|\widehat{\bm{H}}R(\mathcal{S})(\widehat{\bm{G}}\otimes{\bf I}_{K})-(\mathcal{Y}\times_{3}^{1}\bm{X}^{\dagger})^{\langle 1\rangle}\|_{F}^{2},$ (25)
$\min_{\widehat{\bm{G}}\in\mathbb{C}^{N\times UM}}\|({\bf I}_{K}\otimes\widehat{\bm{H}})L(\mathcal{S})\widehat{\bm{G}}-(\mathcal{Y}\times_{3}^{1}\bm{X}^{\dagger})^{\langle 2\rangle}\|_{F}^{2}.$ (26)
Specifically, we apply the following alternating least squares (ALS) algorithm:
$\bm{H}^{(l+1)}=(\mathcal{Y}\times_{3}^{1}\bm{X}^{\dagger})^{\langle 1\rangle}(R(\mathcal{S})(\bm{G}^{(l)}\otimes{\bf I}_{K}))^{\dagger},$
$\bm{G}^{(l+1)}=(({\bf I}_{K}\otimes\bm{H}^{(l+1)})L(\mathcal{S}))^{\dagger}(\mathcal{Y}\times_{3}^{1}\bm{X}^{\dagger})^{\langle 2\rangle}.$
To guarantee the existence of all pseudoinverse matrices, it is imperative that the conditions $N\leq\min\{UMK,LPK\}$ and $T\geq UM$ are met but we note that $N\leq\min\{UMK,LPK\}$ is a sufficient condition, not a necessary one. This is because Equation 12 holds for all coefficients when $T\geq\Omega(UM)$. Additionally, when $\bm{H}$ and $\bm{G}$ exhibit low-rank characteristics, we can employ SVD to truncate small singular values, facilitating the use of low-rank projection in each iteration.
$(ii)$ To address the numerical stability challenges of the pseudoinverse and to accommodate a broader class of pilot signal matrices $\bm{X}$, we can alternatively optimize the following loss function directly:
$\min_{\widehat{\bm{H}}\in\mathbb{C}^{LP\times N}\atop\widehat{\bm{G}}\in\mathbb{C}^{N\times UM}}f(\widehat{\bm{H}},\widehat{\bm{G}})=\frac{1}{T}\|\mathcal{X}([\widehat{\bm{H}},\mathcal{S},\widehat{\bm{G}}])-\mathcal{Y}\|_{F}^{2}.$ (27)
As for (27), we solve the above optimization problem by the following alternating gradient descent (AGD):
$\bm{H}^{(l+1)}=\bm{H}^{(l)}-\mu\nabla_{\bm{H}^{*}}f(\bm{H}^{(l)},\bm{G}^{(l)}),$
$\bm{G}^{(l+1)}=\bm{G}^{(l)}-\mu\nabla_{\bm{G}^{*}}f(\bm{H}^{(l+1)},\bm{G}^{(l)}),$
where $\mu$ is a step size and we define the Wirtinger gradients $\nabla_{\bm{H}^{*}}f(\bm{H}^{(l)},\bm{G}^{(l)})$ and $\nabla_{\bm{G}^{*}}f(\bm{H}^{(l+1)},\bm{G}^{(l)})$ as follows:
[…]
×$\times_{2,3}^{2,3}(\mathcal{S}^{*}\times_{3}^{1}(\bm{G}^{(l)})^{*}),$
[…]
×$\times_{1,2}^{1,2}\big{(}(\mathcal{B}^{(l)}\times_{3}^{1}\bm{X}-\mathcal{Y})\times_{3}^{2}\bm{X}^{*}\big{)}.$
In comparison to the ALS algorithm, the AGD does not require additional coefficient constraints, aside from the condition $T\geq\Omega(UM)$ for satisfying the RIP condition. Additionally, it avoids the numerical stability issues associated with the pseudoinverse.
@@ IV Nonconvex Optimization Algorithms
Figure 1: Recovery performance for channel matrices (a) for different $T$ with $UM=LP=K=N=10$ and $\gamma^{2}=10^{-6}$, (b) for different $UM$ with $T=100$, $LP=K=N=10$ and $\gamma^{2}=10^{-6}$, (c) for different $LP$ with $T=20$, $UM=K=N=10$ and $\gamma^{2}=10^{-6}$.
@@
Figure 2: Recovery performance for channel matrices (a) for different $N$ with $T=20$, $UM=LP=K=10$ and $\gamma^{2}=10^{-6}$, (b) for different $K$ with $T=20$, $UM=LP=N=10$ and $\gamma^{2}=10^{-6}$, $\mu=0.02$ for the $1$ IRS scenario and $\mu=0.1$ for the $2$ IRSs scenario, (c) for different $\gamma^{2}$ with $T=20$ and $UM=LP=K=N=10$.
Multi-hop IRS-assisted MIMO system
In a multi-hop IRS-assisted MIMO system, where channel matrices and IRSs are represented in (20), a closed-form solution for the ALS method is not feasible. Consequently, we primarily consider the AGD method and the following loss function:
[…]
where $\mu$ is a step size and the Wirtinger gradient $\nabla_{\bm{B}_{d}^{*}}g(\bm{B}_{D},\\ \dots,\bm{B}_{0})$ is defined as:
$\nabla_{\bm{B}_{d}^{*}}g(\bm{B}_{D},\dots,\bm{B}_{0})=\frac{1}{T}\sum_{k=1}^{K}\mathcal{S}_{d+1}^{\rm H}(:,k,:)\bm{B}_{d+1}^{\rm H}\cdots$
$\hskip 14.22636pt\mathcal{S}_{D}^{\rm H}(:,k,:)\bm{B}_{D}^{\rm H}\bigg{(}\mathcal{X}([\bm{B}_{D},\mathcal{S}_{D}(:,k,:),\dots,\bm{B}_{1},\mathcal{S}_{1}(:,k,:),$
$\hskip 14.22636pt\bm{B}_{0}])-\widetilde{\mathcal{Y}}(:,k,:)\bigg{)}\bm{X}^{\rm H}\bm{B}_{0}^{\rm H}\cdots\bm{B}_{d-1}^{\rm H}\mathcal{S}_{d}^{\rm H}(:,k,:).$
In line with the single-hop IRS-assisted MIMO system, AGD in this scenario similarly operates without coefficient constraints, except for the condition $T\geq\Omega(UM)$ for satisfying the RIP condition.
We note that for low-rank channel estimation, iterative hard thresholding combined with truncated singular value decomposition [77] can be effectively integrated into the AGD method.
V Experimental Results
In this section, we conduct numerical experiments to evaluate the effectiveness of Theorem 2 and Theorem 5. We consider $1$ IRS and $2$ IRSs assisted MIMO systems, where each element of channel matrices $\bm{H}$, $\bm{G}$, $\bm{B}_{0}$, $\bm{B}_{1}$, $\bm{B}_{2}$, and pilot signal matrix $\bm{X}$ follows independent and identically complex standard normal distribution. Subsequently, we normalize the above channel matrices to have a unit Frobenius norm. To effectively estimate the channel matrices in the $1$ IRS assisted MIMO system, we apply the ALS, AGD, tensor train singular value decomposition (TT-SVD) [37], and tensor train cross-approximation (TT CA) via the greedy restricted cross-interpolation (GRCI) algorithm [78]. As two representative methods for TT decomposition, TT-SVD and TT-CA are employed for the tensor factorization of $\mathcal{Y}\times_{3}^{1}\bm{X}^{\dagger}$. In the case of the $2$ IRSs model, we employ the AGD algorithm. We initialize ALS and AGD with random matrices, with elements generated from a complex standard normal distribution. To simplify parameter selection and notation in the $2$ IRSs system, we set $N_{1}=N_{2}=N$ and denote $\mathcal{B}_{2}=\mathcal{B}$, $\mathcal{B}_{2}^{\star}=\mathcal{B}^{\star}$. The number of iterations for the ALS, AGD ($1$ IRS), and AGD ($2$ IRSs) algorithms are set to $100$, $10^{5}$, and $2\times 10^{5}$, respectively. For the AGD algorithm, we set $\mu=0.2$ for the $1$ IRS scenario and $\mu=1$ for the $2$ IRSs scenario. For each experimental setting, we conduct 20 Monte Carlo trials and then take the average over the 20 trials to report the results.
In the first experiment, we compare the recovery performance of different methods with various values of $T$, $UM$, and $LP$ in Figure 1. These plots illustrate that the recovery error tends to decrease with larger values of $T$, or smaller values of $UM$ and $LP$, consistent with our theoretical findings presented in Theorem 2 and Theorem 5. Additionally, we observe the following: $(i)$ when $T$ approaches $UM$, the recovery error could worsen since RIP conditions (9) and (21) are not satisfied; $(ii)$ AGD (1 IRS) and ALS exhibit similar recovery performance except for when $T=UM$; $(iii)$ the recovery error of AGD (2 IRSs) is larger than AGD (1 IRS) due to the introduction of an additional channel matrix, leading to more error as shown in Theorem 5; $(iv)$ compared to the ALS and AGD algorithms, TT-SVD and TT CA demonstrate poorer recovery performance. This is because TT-SVD and TT CA do not utilize the known information of $\mathcal{S}$ or $\mathcal{S}_{i},i=1,2$, resulting in a larger error. Furthermore, the recovery error of TT-CA is larger than that of TT-SVD, as TT-SVD utilizes the full information of the channel tensor, whereas TT-CA does not [79]; $(v)$ from Figure 1, it is evident that successful recovery is not dependent on $LP$, as even with $LP=100$, the recovery remains stable when $T=20$.
In the second experiment, we examine the recovery performance of different methods across varying values of $N$ and $K$. Keeping a fixed number of time slots $T=20$, we observe stable recovery errors with increasing $N$ and $K$. However, it is noteworthy that as $K$ increases, the recovery performance generally improves, except for TT-SVD and TT CA. This is attributed to the fact that the coherence time $T_{s}=KT$ becomes larger as $K$ increases while $T$ remains fixed, thereby enhancing performance. However, factorization approaches like TT-SVD and TT CA struggle to effectively denoise as $K$ increases. This is due to the possibility that the estimated $\widehat{\mathcal{S}}$ or $\widehat{\mathcal{S}}_{1},\widehat{\mathcal{S}}_{2}$ may introduce more noise under these conditions. This highlights a limitation of tensor factorization in recovering channel matrices of IRS-based systems.
In the third experiment, we compare the recovery performance of different methods across various noise variances $\gamma^{2}$. In Figure 2, we observe that, except for TT CA, the recovery error exhibits a linear relationship with $\gamma^{2}$, consistent with the analysis in Theorem 2 and Theorem 5. Ultimately, while the ALS achieves performance similar to the AGD, the AGD may offer greater stability, especially in time-varying systems where the pseudoinverse of the input matrix may not be well-defined.
@@ V Experimental Results
Figure 3: Recovery performance for channel matrices for different $D$ with $K=N_{0}=\cdots N_{D}=10$, $\gamma^{2}=10^{-6}$ and $\mu$ sequentially selected from the set $\{0.2,1,5,10,15\}$.
In the final experiment, we compare the recovery performance of AGD for different numbers of IRSs $D$. As shown in Figure 3, the recovery error increases linearly with $D$, aligning with the theoretical analysis in Theorem 5. Additionally, as $T$ increases, the recovery error decreases. Thus, for fixed $T$ time slots, a trade-off emerges between mitigating propagation attenuations with multiple IRSs and minimizing recovery error.
VI Conclusion
@@
This paper addresses the significant gap in theoretical error analysis concerning channel estimation for IRS-assisted MIMO systems. By establishing the equivalence between these systems and tensor train-based tensor-on-tensor (ToT) regression, we provide insights into the fundamental factors crucial for the successful recovery of channel matrices. Our analysis highlights the pivotal role of the relationship between the number of user terminals or base stations and the number of time slots for ensuring stable recovery. Furthermore, we extend our investigation to consider low-rank channel matrices and unknown IRS, enhancing the applicability of our findings. Additionally, through our exploration of a multi-hop IRS scheme, we evaluate corresponding recovery errors, shedding light on the performance of such configurations. Finally, to validate our theoretical conclusions, we propose and implement two nonconvex optimization algorithms–alternating least squares and alternating gradient descent–demonstrating their effectiveness through experimental results.
@@
An important area for future research involves conducting error analysis on two-way, two-hop IRS-assisted MIMO systems [80, 81, 82]. Unlike the one-way multi-hop IRS-assisted model discussed in this paper, the optimization problem in the two-way scenario, where channels between base stations/user terminals and IRSs are accessible, consists of two ToT regression problems. Thus, our current analytical approach cannot be directly extended to this setup. Furthermore, another challenging aspect for future investigation is the local convergence analysis and global geometry of the IRS-assisted optimization problem. Despite the promising experimental results obtained even with random initialization, unlike the rotation ambiguity commonly encountered in standard nonconvex matrix/tensor recovery [45, 46, 47, 48], a diagonal (scaling) ambiguity exists between channel matrices, as noted in [36, Section 2]. This ambiguity poses a challenge in analyzing the optimization problem at hand.

### v2

@@ III-C Discussion
IV Experimental Results
In this section, we present numerical experiments to validate the effectiveness of the bounds in Theorems 2, 3, 5 and 6 for single-hop and multi-hop IRS-assisted MIMO systems [^5] Here, we replace $\delta_{UM}$ with $c\sqrt{\frac{UM}{T}}$, where $c$ is a positive constant, in accordance with (10) and (26). Moreover, since the performance metric in these experiments is the normalized recovery error $\frac{\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}^{2}}{\|\mathcal{B}^{\star}\|_{F}^{2}}$, we accordingly normalize the upper and lower bounds in Theorems 2, 3, 5 and 6 by dividing them by $\|\mathcal{B}^{\star}\|_{F}^{2}$.. To simplify parameter selection, each entry of the channel matrices and the pilot signal matrix is independently drawn from the standard complex normal distribution. The channel matrices are then normalized to have unit Frobenius norm. Additionally, we set $N_{1}=\cdots=N_{D-1}=N$ and denote $\mathcal{B}_{D}=\mathcal{B}$ and $\mathcal{B}_{D}^{\star}=\mathcal{B}^{\star}$ for notational simplicity. To effectively estimate the channel matrices in the single-hop IRS-assisted MIMO system, we adopt the bilinear alternating least squares (BALS) algorithm [31], which is designed for constrained least squares minimization and has been widely used as a benchmark in channel estimation [32, 33, 84, 85, 86]. In contrast to BALS, which admits closed-form updates in the single-hop case, the constrained least squares problem in the multi-hop IRS-assisted MIMO system ($D\geq 2$) lacks an analytical solution. To address this, we adopt the alternating gradient descent (AGD) algorithm with learning rate $\mu$, following the gradient-based strategy proposed in [77]. The implementation details and computational complexities of both algorithms are provided in the Appendix B. We initialize BALS and AGD with random matrices, with elements generated from a complex standard normal distribution. The number of iterations is set to $100$ for BALS and $10000$ for AGD. For each experimental setting, we perform 20 Monte Carlo trials and report the average results across these trials.
In the first experiment, we investigate the recovery performance of the BALS algorithm under varying values of $UM$ and $LP$, as depicted in Figure 1. The result illustrates that the recovery error consistently falls within the theoretical range defined by the upper and lower bounds in Theorems 2 and 3. In particular, the observed recovery error closely matches the upper bound associated with constrained least squares minimization, which scales proportionally with $LPN+UMN$, as characterized in Theorem 2. This scaling behavior reflects the intrinsic nonconvex structure of the IRS-assisted MIMO channel estimation problem, where $LPN$ and $UMN$ correspond to the degrees of freedom for the two channel matrices involved in the estimation problem. The empirical agreement with the theoretical bound further confirms the tightness and optimality of our recovery guarantees.
@@ IV Experimental Results
Figure 1: Recovery performance of channel matrices in the single-hop IRS assisted MIMO system for different UMUM and $LP$ with $K=N=10$, $T=100$ and $\gamma^{2}=10^{-6}$.
In the second experiment, we systematically examine the recovery performance of the BALS algorithm across varying values of $N$, $T$ and $\gamma^{2}$. The empirical findings illustrated in Figure 2 demonstrate strong concordance with our theoretical guarantees. Notably, the recovery error exhibits an increasing trend with larger $N$ and $\gamma^{2}$, and a decreasing trend as the number of time slots $T$ grows, reflecting the fundamental trade-offs inherent in the estimation problem. Crucially, the observed recovery errors are rigorously confined within the analytically derived upper and lower bounds specified in Theorems 2 and 3, thereby substantiating the tightness and validity of our theoretical framework.
@@
Figure 3: Recovery performance of channel matrices in the single-hop IRS assisted MIMO system for different $\gamma_{1}^{2}$ with $UM=LP=K=N=10$, $\gamma^{2}=10^{-6}$ and $T=100$.
In the preceding analysis, the BS-UT direct channels were omitted when evaluating the recovery performance in the single-hop IRS-assisted MIMO system. However, such direct links are often present in practical deployments. To account for this, the model in (2) can be reformulated as $\bm{Y}(k)=\bm{H}\text{diag}(\bm{s}(k))\bm{G}\bm{X}+\bm{A}\bm{X}+\bm{W}(k),k\in[K]$, where $\bm{A}\in\mathbb{C}^{LP\times UM}$ represents the BS-UT channel matrix. In the third experiment, we investigate the impact of the BS-UT channel $\bm{A}$ on the recovery performance of the BALS algorithm. Importantly, the recovery objective remains focused on estimating $\bm{H}$ and $\bm{G}$, rather than jointly estimating $\bm{H}$, $\bm{G}$, and $\bm{A}$. We further assume that each entry of $\bm{A}$ follows $\mathcal{C}\mathcal{N}(0,\gamma_{1}^{2})$. As shown in Figure 3, the recovery error increases as $\gamma_{1}^{2}$ becomes larger. This behavior arises because the term $\bm{A}\bm{X}$ effectively acts as an additional noise component when optimizing $\bm{H}$ and $\bm{G}$, similar in effect to $\bm{W}(k)$. Specifically, the $(i,j)$-th entry of $\bm{A}\bm{X}+\bm{W}(k)$ follows a complex Gaussian distribution $\mathcal{C}\mathcal{N}(0,\gamma_{1}^{2}\|\bm{X}(:,j)\|_{2}^{2}+\gamma^{2})$. As a result, the conclusions in Theorems 2 and 3 can be naturally extended by replacing $\gamma^{2}$ with $\max_{j}\gamma_{1}^{2}\|\bm{X}(:,j)\|_{2}^{2}+\gamma^{2}$. The empirical recovery errors remain tightly bounded by the theoretically derived upper and lower limits, thus confirming the validity of the theoretical analysis.
In the fourth experiment, we evaluate the recovery performance of the AGD algorithm in multi-hop IRS-assisted MIMO systems with $D=2$, under varying values of $UM$, LPLP, NN, $T$ and $\gamma^{2}$. As shown in Figures 4, 4, 4, 5 and 5, the observed recovery errors exhibit trends that are consistent with the theoretical predictions in Theorems 5 and 6. Moreover, the empirical recovery errors remain tightly confined within the theoretically established upper and lower bounds, thereby providing additional evidence for the optimality of the derived recovery guarantees.
@@
Figure 4: Recovery performance of channel matrices in the multi-hop IRS assisted MIMO system (a) for different UMUM with $D=2$, $\mu=1$, $LP=K=N=10$, $T=100$ and $\gamma^{2}=10^{-6}$, (b) for different LPLP with $D=2$, $\mu=1$, $UM=K=N=10$, $T=100$ and $\gamma^{2}=10^{-6}$, (c) for different NN with $D=2$, $\mu=1$, $T=UM=LP=K=10$ and $\gamma^{2}=10^{-6}$.
Figure 5: Recovery performance of channel matrices in the multi-hop IRS assisted MIMO system (a) for different $T$ with $D=2$, $\mu=1$, $UM=LP=K=N=10$ and $\gamma^{2}=10^{-6}$, (b) for different $\gamma^{2}$ with $D=2$, $\mu=1$, $UM=LP=K=N=10$ and $T=20$, (c) for different $D$ with $T=UM=LP=K=N=10$, $\gamma^{2}=10^{-6}$ and $\mu$ sequentially selected from the set $\{1,5,10,15\}$.
In the final experiment, we compare the recovery performance of AGD for different numbers of IRSs $D$. As shown in Figure 5, the recovery error increases linearly with $D$, aligning with the theoretical analysis in Theorems 5 and 6. Moreover, for a fixed number of time slots $T$, the results highlight a fundamental trade-off between enhancing signal propagation through additional IRS layers and maintaining low recovery error. Note that while increasing $D$ can improve end-to-end channel coverage, it simultaneously escalates the difficulty of accurate channel estimation due to the higher dimensionality and the accumulation of noise.
V Conclusion
@@
This paper addresses the significant gap in optimal error analysis concerning channel estimation for IRS-assisted MIMO systems. By establishing the equivalence between these systems and tensor train-based ToT regression, we provide insights into the fundamental factors crucial for the successful recovery of channel matrices. Specifically, we have established a theoretical upper bound on the recovery error for the constrained least-squares solution, along with a minimax lower bound for the IRS-assisted MIMO model. Our results demonstrate that the recovery error decreases inversely with the number of time slots and increases proportionally with the number of unknown entries in the channel matrices. These findings quantify the fundamental trade-offs inherent in IRS-assisted MIMO channel estimation and provide theoretical justification for the observed empirical behaviors. Additionally, through our exploration of a multi-hop IRS scheme, we evaluate corresponding recovery errors, shedding light on the performance of such configurations. Finally, we validate our theoretical findings through extensive simulations, which consistently align with the derived recovery bounds.
Building upon the results established in this paper, several important research directions can be identified for further investigation.
@@
  • An important area for future research involves conducting error analysis on two-way, two-hop IRS-assisted MIMO systems [87, 88, 89]. Unlike the one-way multi-hop IRS-assisted model discussed in this paper, the optimization problem in the two-way scenario, where channels between base stations/user terminals and IRSs are accessible, consists of two ToT regression problems. Thus, our current analytical approach cannot be directly extended to this setup.
  • Another important challenge is extending our results to the downlink communication scenario. Notably, our recovery guarantees remain applicable in the downlink setting, provided that the RIP holds–achieved by simply interchanging the roles of transmitter and receiver. However, as evidenced by prior studies [90, 91, 92, 93], the number of time slots in the downlink can fall below the number of BSs, thereby violating the RIP condition. This limitation underscores the necessity of developing new analytical frameworks tailored to the downlink regime.
  • Furthermore, recent studies have revealed that even when the direct BS–UT link is blocked, the scattering matrix of the direct channel remains nonzero [94]. Under specific assumptions, the IRS-assisted MIMO model can therefore be reformulated as in [95, eq. (80)], which explicitly incorporates the scattering matrix of the reconfigurable impedance network and the IRS reflection (or scattering) coefficient matrix. When these matrices are known, our theoretical results remain directly applicable. However, in practical scenarios, the scattering matrix is typically unknown and the model includes an inverse operation, which prevents the direct application of our current analytical approach. Consequently, extending the error analysis to models without explicit scattering information remains an open and important direction for future work.
  • Finally, a further avenue for exploration is the local convergence analysis and global geometry of the IRS-assisted optimization problem. Despite the promising experimental results obtained even with random initialization, unlike the rotation ambiguity commonly encountered in standard nonconvex matrix/tensor recovery [47, 48, 49, 51], a diagonal (scaling) ambiguity exists between channel matrices, as noted in [37, Section 2]. This ambiguity poses a challenge in analyzing the optimization problem at hand.